# Confidence on the Focal: Conformal Prediction with Selection-Conditional Coverage

### v1

@@ 2 Problem setup
We are interested in inferring the unknown outcome $Y$ of certain “focal” units selected in a data-driven manner. Following the split conformal prediction framework (Vovk et al.,, 2005; Lei et al.,, 2018), our prediction set is based on a fitted prediction model obtained from a training set ${\mathcal{D}}_{\textnormal{train}}$. Recall the calibration data ${\mathcal{D}}_{\textnormal{calib}}=\{Z_{i}\}_{i=1}^{n}$, $Z_{i}=(X_{i},Y_{i})$, and test points $\{Z_{n+j}\}_{j=1}^{m}$, $Z_{n+j}=(X_{n+j},Y_{n+j})$. The outcomes $\{Y_{n+j}\}_{j=1}^{m}$ are unobserved and of interest. We write ${\mathcal{D}}_{\textnormal{test}}=\{X_{n+j}\}_{j=1}^{m}$ for brevity. We assume ${\mathcal{D}}_{\textnormal{train}}$ is independent of ${\mathcal{D}}_{\textnormal{calib}}\cup{\mathcal{D}}_{\textnormal{test}}$, and we condition on ${\mathcal{D}}_{\textnormal{train}}$, treating the fitted models as fixed.
@@ 2.1 Selection-conditional coverage
We call $Z_{n+j}$ a focal unit if $j\in\widehat{\mathcal{S}}$, where $\widehat{\mathcal{S}}={\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})\subseteq[m]$ is obtained from a selection rule ${\mathcal{S}}$ that may depend on the observe data. Its potential dependence on ${\mathcal{D}}_{\textnormal{train}}$ is clear since we treat ${\mathcal{D}}_{\textnormal{train}}$ as fixed. Without loss of generality, we posit that ${\mathcal{S}}$ is a determinististic function; when the selection rule is randomized, one can condition on its randomness and follow our framework.
@@
where $\alpha\in(0,1)$ is the confidence level, and $\mathfrak{S}\subseteq 2^{[m]}$ is some pre-specified collection of subsets of $[m]$. We call $\mathfrak{S}$ the selection taxonomy in what follows.
Equation (3) is the most general form of the selection-conditional coverage we aim for in this paper, where difference choices of $\mathfrak{S}$ leads to a spectrum of granularity in the conditioning event. For instance, taking $\mathfrak{S}=\{S_{0}\}$ for some $S_{0}\subseteq[m]$ achieves coverage conditional on selecting a specific set $S_{0}$; this is similar to the coverage guarantee in (Lee et al.,, 2016) for high-dimensional inference. Taking $\mathfrak{S}=\{S\subseteq[m]\colon|S|=r\}$ for some $r\leq m$ achieves coverage conditional on getting a specific number of selected units. Finally, taking $\mathfrak{S}=2^{[m]}$ puts no restrictions on the selection set, giving rise to the guarantee (2) introduced in the beginning, coverage conditional only on a unit being selected.
@@ 2.2 Relations among notions of selective coverage
Before introducing our methods, we take a moment here to compare different notions of selective coverage. Readers who are more interested in our methodology may skip the remaining of this section.
Our first observation is that (3) implies (2) under appropriate conditions. The proof of the next Proposition is in Appendix B.1.
@@
If $\widehat{C}_{\alpha,n+j}$ satisfies (3) for a set of mutually disjoint taxonomies $\{\mathfrak{S}_{\ell}\}_{\ell\in{\mathcal{L}}}$ such that $\cup_{\ell\in{\mathcal{L}}}\mathfrak{S}_{\ell}=2^{[m]}$, then it satisfies (2).
To distinguish the selection-conditional coverage in (3) and that in (2), we refer to the former as strong selection-conditional coverage and the latter as weak selection-conditional coverage.
Another widely used post-selection-type guarantee is the false coverage rate (FCR) (Benjamini and Yekutieli,, 2005), defined as the expected proportion of selected units missed by the prediction set:
@@
where $a\vee b=\max\{a,b\}$ for any $a,b\in\mathbb{R}$. As introduced earlier, Bao et al., (2024) mainly focuses on constructing prediction sets with FCR control. However, with $m=1$, it holds that $\text{FCR}=\mathbb{P}(j\in\widehat{{\mathcal{S}}},Y_{n+j}\notin\widehat{C}_{\alpha}(X_{n+j}))\leq\alpha$ for any marginally valid prediction set. Thus, the FCR is not very meaningful when we consider one test point.
We claim that the strong selection-conditional coverage implies FCR control with a proper choice of the selection taxonomy. The proof of the following proposition can be found in Appendix B.2.
@@
[…]
As a side note, we find that the weak selection-conditional coverage does not necessarily imply FCR control, although in some special cases both can be true (see, e.g., Bao et al., (2024) for such examples). We put this as a proposition below, with a counterexample given in Appendix B.3.
@@
There exists an instance and prediction sets $\widehat{C}_{\alpha}$ such that $\widehat{C}_{\alpha}$ satisfies the weak selection-conditional coverage at level $\alpha$ but violates the FCR control at level $\alpha$.
We close this part by introducing several notations that will be used throughout the paper.
Notations.
We denote the index sets ${\mathcal{I}}_{{\textnormal{calib}}}=\{1,\dots,n\}$ and ${\mathcal{I}}_{\textnormal{test}}=\{n+1,\dots,n+m\}$. Let $S:{\mathcal{X}}\times{\mathcal{Y}}\mapsto\mathbb{R}$ be the score function for selection, and $V:{\mathcal{X}}\times{\mathcal{Y}}\mapsto\mathbb{R}$ the nonconformity score function for constructing the prediction set. For example, $S(x,y)=y$ and $V(x,y)=|y-\widehat{\mu}(x)|$, where $\widehat{\mu}(x)$ is regression function fitted on ${\mathcal{D}}_{\textnormal{train}}$. For any $j\in{\mathcal{D}}_{\textnormal{calib}}\cup{\mathcal{D}}_{\textnormal{test}}$, define $S_{i}=S(X_{i},Y_{i})$ and $V_{j}=V(X_{j},Y_{j})$.
Let ${\mathcal{D}}_{j}={\mathcal{D}}_{\textnormal{calib}}\cup\{Z_{n+j}\}$, and ${\mathcal{I}}_{j}={\mathcal{I}}_{\textnormal{calib}}\cup\{n+j\}$ be the corresponding indices; define also ${\mathcal{D}}^{c}_{j}={\mathcal{D}}_{\textnormal{test}}\backslash\{Z_{n+j}\}$ and ${\mathcal{I}}^{c}_{j}={\mathcal{I}}_{\textnormal{test}}\backslash\{n+j\}$. For any $j\in[m]$, we define $Z_{n+j}(y)=(X_{n+j},y)$, the value of $Z_{n+j}$ if $Y_{n+j}$ were to take a hypothesized value $y\in{\mathcal{Y}}$.
@@ 3.1 Warm-up: split conformal prediction
To warm up, we briefly summarize the split conformal prediction (SCP) method (Vovk et al.,, 2005; Lei et al.,, 2018) and the ideas for its finite-sample coverage due to exchangeability.
We begin with a nonconformity score function $V\colon{\mathcal{X}}\times{\mathcal{Y}}\mapsto\mathbb{R}$ determined by the training set, so that $V(x,y)$ informs how well a hypothetical value $y\in{\mathcal{Y}}$ conforms to a machine prediction. For instance, one may set $V(x,y)=|y-\widehat{\mu}(x)|$, where $\widehat{\mu}(x)$ is regression function fitted on ${\mathcal{D}}_{\textnormal{train}}$. Other popular choices include CQR (Romano et al.,, 2019) for regression and APS (Romano et al.,, 2020) for classification. We compute $V_{i}=V(X_{i},Y_{i})$ for $i\in[n]$. The split conformal prediction set for a test point $j\in[m]$ is then
@@
where $\text{Quantile}(1-\alpha;\cdot)$ is the empirical $(1-\alpha)$-th quantile of the set in the second argument. When $(Z_{1},\dots,Z_{n},Z_{n+j})$ are exchangeable, $\widehat{C}_{\alpha,n+j}^{\text{SCP}}$ achieves valid marginal coverage (1(Vovk et al.,, 2005).
It helps motivate our approach to first introduce the ideas behind the validity of $\widehat{C}_{\alpha,n+j}^{\text{SCP}}$. Note that
@@
where $V_{n+j}=V(X_{n+j},Y_{n+j})$. Define the unordered set $[Z]_{j}=[Z_{1},\dots,Z_{n},Z_{n+j}]$. Due to exchangeability, conditional on the event $\{[Z]_{j}=[z_{1},\dots,z_{n+j}]\}$, the probability of $Z_{n+j}$ taking on each value in $z_{1},\dots,z_{n+j}$ is equal. The right-handed side of (6) is thus lower bounded by $1-\alpha$.
The key challenge of achieving selection-conditional coverage like (3) is that, conditional on the selection event, the data points $\{Z_{1},\dots,Z_{n},Z_{n+j}\}$ are not exchangeable any more. More specifically, given selection, and further conditional on $\{[Z]_{j}=[z_{1},\dots,z_{n+j}]\}$, it is unclear how $Z_{n+j}$ is distributed over $z_{1},\dots,z_{n+j}$. As such, a natural idea is to find a subset of calibration data that are “exchangeable” with the new test point even given the selection event. We introduce our methods in the next.
3.2 Conformal inference via locally exchangeable subgroups
Fix a test unit $j\in[m]$. Recall that $\widehat{\mathcal{S}}={\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})$ is the focal set, and $Z_{n+j}(y)=(X_{n+j},y)$ is the imputed data point with a hypothesized response $y\in{\mathcal{Y}}$. For any calibration data $i\in{\mathcal{I}}_{\textnormal{calib}}$, we define the “swapped” calibration data ${\mathcal{D}}_{{\textnormal{calib}}}^{{\textnormal{swap}(i,j)}}(y)$ and test data ${\mathcal{D}}_{{\textnormal{test}}}^{{\textnormal{swap}(i,j)}}(y)$ to be

### v3

@@ 2 Problem setup
Following the split conformal prediction framework (Vovk et al.,, 2005; Lei et al.,, 2018), we build our prediction sets based on a prediction model fitted on a training set ${\mathcal{D}}_{\textnormal{train}}$, assuming ${\mathcal{D}}_{\textnormal{train}}$ is independent of the calibration and test data. In what follows, we always condition on ${\mathcal{D}}_{\textnormal{train}}$, thereby treating the fitted models as fixed. In this section, we formally introduce the selection-conditional coverage guarantees and compare them to other related notions in the literature.
@@ 2.1 Selection-conditional coverage
Recall that $j\in[m]$ is a focal unit if $j\in\widehat{\mathcal{S}}$, where $\widehat{\mathcal{S}}={\mathcal{S}}({\mathcal{D}}_{\textnormal{calib}},{\mathcal{D}}_{\textnormal{test}})\subseteq[m]$ is obtained from a selection rule ${\mathcal{S}}$ that depends on the observed data. Its potential dependence on ${\mathcal{D}}_{\textnormal{train}}$ is clear since we treat ${\mathcal{D}}_{\textnormal{train}}$ as fixed. Without loss of generality, we posit that ${\mathcal{S}}$ is a deterministic function; when the selection rule is randomized, one can condition on its randomness and follow our framework.
@@
where $\alpha\in(0,1)$ is the confidence level, and $\mathfrak{S}\subseteq 2^{[m]}$ is some pre-specified collection of subsets of $[m]$. We call $\mathfrak{S}$ the selection taxonomy in what follows. Different choices of $\mathfrak{S}$ lead to a spectrum of granularity in the conditioning event. For instance, taking $\mathfrak{S}=2^{[m]}$ puts no restrictions on the selection set, giving rise to the guarantee (2) introduced in the beginning (coverage conditional on a unit being selected). Taking $\mathfrak{S}=\{S\subseteq[m]\colon|S|=r\}$ for some $r\leq m$ achieves coverage conditional on selecting a specific number of units. Finally, taking $\mathfrak{S}=\{S_{0}\}$ for some $S_{0}\subseteq[m]$ achieves coverage conditional on selecting a specific set; this is similar to the coverage guarantee conditional on a selected model in Lee et al., (2016) for high-dimensional parameter inference.
In words, the selection-conditional coverage guarantee (3) can be interpreted as follows: imagine there are infinitely many independent realizations of ${\mathcal{D}}_{\textnormal{calib}}$ and ${\mathcal{D}}_{\textnormal{test}}$; among those realizations where the selection event of interest happens, the prediction set $\widehat{C}_{\alpha,n+j}$ will cover the true outcome for at least $1-\alpha$ fraction of times. Here, $\mathfrak{S}$ is introduced to provide practitioners with the language to specify the granularity of the conditioning event, allowing them to tailor the guarantee to their specific needs. Moreover, it allows us to properly describe the relationship between selection-conditional coverage and FCR control, as we will see below.
@@ 2.2 Relations among notions of selective coverage
Before introducing our methods, we take a moment to compare different notions of selective coverage. Readers who are more interested in our methodology may skip the remaining of this section.
Our first observation is that (3) implies (2) under appropriate conditions. The proof of the next proposition is in Supplementary Section D.1.
@@
Suppose a family of prediction sets $\{\widehat{C}_{\alpha,n+j}^{(\ell)}\}_{\ell\in{\mathcal{L}}}$ satisfy $\mathbb{P}(Y_{n+j}\in\widehat{C}_{\alpha,n+j}^{(\ell)}{\,|\,}j\in\widehat{\mathcal{S}},\leavevmode\nobreak\ \widehat{\mathcal{S}}\in\mathfrak{S}_{\ell})\geq 1-\alpha$ for a set of disjoint taxonomies $\{\mathfrak{S}_{\ell}\}_{\ell\in{\mathcal{L}}}$ such that $\cup_{\ell\in{\mathcal{L}}}\mathfrak{S}_{\ell}=2^{[m]}$. Define the prediction set $\widehat{C}_{\alpha,n+j}=\widehat{C}_{\alpha,n+j}^{(\ell)}$ when $j\in\widehat{\mathcal{S}}$ and $\widehat{\mathcal{S}}\in\mathfrak{S}_{\ell}$. Then $\mathbb{P}(Y_{n+j}\in\widehat{C}_{\alpha,n+j}{\,|\,}j\in\widehat{\mathcal{S}})\geq 1-\alpha$.
To distinguish the selection-conditional coverage in (3) and that in (2), we refer to (3) as strong selection-conditional coverage and (2) as weak selection-conditional coverage.
Another widely used post-selection guarantee is the false coverage rate (FCR) (Benjamini and Yekutieli,, 2005), defined as the expected proportion of selected units missed by the prediction set:
@@
where $a\vee b=\max\{a,b\}$ for any $a,b\in\mathbb{R}$. Previous works (Weinstein and Ramdas,, 2020; Bao et al.,, 2024; Gazin et al.,, 2024) mainly focus on constructing prediction sets with FCR control. However, with $m=1$, it holds that $\text{FCR}=\mathbb{P}(j\in\widehat{{\mathcal{S}}},Y_{n+j}\notin\widehat{C}_{\alpha}(X_{n+j}))\leq\alpha$ for any marginally valid prediction set. Thus, FCR does not always address the selection issue.
The following proposition shows that strong selection-conditional coverage implies FCR control with a proper choice of selection taxonomy, whose proof is in Supplementary Section D.2.
@@
Suppose a family of prediction sets $\{\widehat{C}_{\alpha,n+j}^{(\ell)}\}_{\ell\in{\mathcal{L}}}$ satisfy $\mathbb{P}(Y_{n+j}\in\widehat{C}_{\alpha,n+j}^{(\ell)}{\,|\,}j\in\widehat{\mathcal{S}},\leavevmode\nobreak\ \widehat{\mathcal{S}}\in\mathfrak{S}_{\ell})\geq 1-\alpha$ for a set of disjoint selection taxonomies $\{\mathfrak{S}_{\ell}\}_{\ell\in{\mathcal{L}}}$ such that $\cup_{\ell\in{\mathcal{L}}}\mathfrak{S}_{\ell}=2^{[m]}$, and for each $\ell\in{\mathcal{L}}$, $\mathfrak{S}_{\ell}\subseteq\{S\subseteq[m]\colon|S|=r(\ell)\}$ for some $0\leq r(\ell)\leq m$. Define the prediction set $\widehat{C}_{\alpha,n+j}=\widehat{C}_{\alpha,n+j}^{(\ell)}$ when $j\in\widehat{{\mathcal{S}}}$ and $\widehat{\mathcal{S}}\in\mathfrak{S}_{\ell}$. Then its FCR (4) is upper bounded by $\alpha\cdot\mathbb{P}(\widehat{{\mathcal{S}}}\neq\varnothing)\leq\alpha$.
The selection taxonomies in Proposition 2 require the selection set to be of a specific size, without other conditions on its form. This can be automatically satisfied by some selection rules such as top-K selection. Extending Proposition 2 to non-exact scenarios yields useful insights in practice. For instance, consider the family of taxonomies with $\mathfrak{S}_{\ell}\subseteq\{S\subseteq[m]\colon r(\ell)-\delta\leq|S|\leq\delta+r(\ell)\}$ for some small $\delta\in\mathbb{N}^{+}$. When the selection size is stable, weak selection-conditional coverage implies $\mathbb{P}(Y_{n+j}\in\widehat{C}_{\alpha,n+j}{\,|\,}j\in\widehat{\mathcal{S}},\leavevmode\nobreak\ \widehat{\mathcal{S}}\in\mathfrak{S}_{\ell})\approx 1-\alpha$, which leads to $\textrm{FCR}\approx\alpha\cdot\mathbb{P}(\widehat{{\mathcal{S}}}\neq\varnothing)$. This approximate result helps explain several phenomena in our numerical experiments: first, the FCR is usually controlled empirically by our method; second, we sometimes observe significantly lower FCR than the selection-conditional coverage when $\mathbb{P}(\widehat{{\mathcal{S}}}=\varnothing)$ is moderately large. As we shall see later, a version of our proposed method achieves coverage guarantees conditional on being selected and the selection size, which requires checking a slightly more complex condition in the reference set construction when the selection size varies. The computation of such prediction sets can be done efficiently for all the instances provided in the paper, with the corresponding worst-case computational complexity explicitly stated in each section. One potential concern, though, is that the additional condition on the selection size may reduce the number of calibration data points in the reference set, especially when the selection step is highly variable. This could potentially lead to wider and/or unstable prediction intervals.
As a side note, the weak selection-conditional coverage does not necessarily imply FCR control, although in some special cases both can be true (see, e.g., Bao et al., (2024) for such examples). We put this as a proposition below, with a counterexample given in Supplementary Section D.3.
@@
There exists an instance and prediction sets $\{\widehat{C}_{\alpha,n+j}:j\in\widehat{{\mathcal{S}}}\}$ that satisfy the weak selection-conditional coverage at level $\alpha$ but violate the FCR control at level $\alpha$.
We end this section with a remark on the interpretation of selection-conditional coverage and FCR control.
Remark 1 (Interpretation of selection-conditional coverage and FCR control).
As shown by Proposition 2, the guarantee in (3) is stronger than FCR control (for a proper choice of the selection taxonomy). In fact, the former has often been used as a device to derive the latter in the literature (e.g., Weinstein et al.,, 2013; Bao et al.,, 2024).
In terms of interpretation, there are two main differences between selection-conditional coverage and FCR. First, the guarantee of FCR is averaged over all the selection events, including the case of empty selection set where the false coverage proportion is by definition zero. Therefore, even with FCR control, one could still suffer from a high false coverage proportion when the selection set is non-empty as long as this is compensated by the cases where the selection set is empty. On the other hand, selection-conditional coverage delivers guarantees conditioning on selection events of interest, which prevents the aforementioned undesired situation. Second, FCR control provides a guarantee that is averaged over all the selected units: it could be possible that for some selected units, the coverage is much lower than the nominal level, and for others the coverage is much higher, so that the average coverage over all the selected units is controlled at the nominal level. In contrast, selection-conditional coverage provides guarantees specific to each selected units.
Finally, we note that the distinction between the two concepts can be asymptotically negligible in special cases. We refer to Lemma 1 in the supplementary material for such an instance.
@@ 3.1 Warm-up: split conformal prediction
To warm up, we briefly summarize the split conformal prediction (SCP) method (Vovk et al.,, 2005; Lei et al.,, 2018) and how it achieves finite-sample coverage under exchangeability.
SCP starts with a nonconformity score function $V\colon{\mathcal{X}}\times{\mathcal{Y}}\mapsto\mathbb{R}$ determined by ${\mathcal{D}}_{\textnormal{train}}$, so that $V(x,y)$ informs how well a hypothetical value $y\in{\mathcal{Y}}$ conforms to a machine prediction. For instance, one may set $V(x,y)=|y-\widehat{\mu}(x)|$, where $\widehat{\mu}(x)$ is regression function fitted on ${\mathcal{D}}_{\textnormal{train}}$. Other popular choices include conformalized quantile regression (CQR, Romano et al.,, 2019) for regression and adaptive prediction sets (APS, Romano et al.,, 2020) for classification.
Compute $V_{i}=V(X_{i},Y_{i})$ for $i\in[n]$. The split conformal prediction set for test unit $j\in[m]$ is
@@
where $\text{Quantile}(1-\alpha;\cdot)$ is the $(1-\alpha)$-th empirical quantile of the set in the second argument. When $(Z_{1},\dots,Z_{n},Z_{n+j})$ are exchangeable, $\widehat{C}_{\alpha,n+j}^{\text{SCP}}$ achieves (1(Vovk et al.,, 2005).
It helps motivate our approach to see the ideas behind the validity of $\widehat{C}_{\alpha,n+j}^{\text{SCP}}$. In words, $\widehat{C}_{\alpha,n+j}^{\text{SCP}}$ finds hypothesized values of $y$ that make $V(X_{n+j},y)$ look similar to calibration scores. Note that
@@
where $V_{n+j}=V(X_{n+j},Y_{n+j})$. Recall the unordered set $[{\mathcal{D}}_{j}]=[Z_{1},\dots,Z_{n},Z_{n+j}]$ where $Z_{i}=(X_{i},Y_{i})$. Conditional on the event $\{[{\mathcal{D}}_{j}]=[z_{1},\dots,z_{n+j}]\}$, the only randomness is in the ordering of $(Z_{1},\dots,Z_{n},Z_{n+j})$; due to exchangeability, the probability of $Z_{n+j}$ taking on each value in $z_{1},\dots,z_{n},z_{n+j}$ is equal. Therefore, conditional on $[{\mathcal{D}}_{j}]=[z_{1},\dots,z_{n+j}]$, the chance of $V_{n+j}$ being no greater than the $(1-\alpha)$-th quantile of $[v_{1},\dots,v_{n},v_{n+j}]$, where $v_{i}=v(z_{i})$, is at least $1-\alpha$, i.e.,
$\mathbb{P}\Big{(}V_{n+j}\leq\text{Quantile}\big{(}1-\alpha;\{v_{i}\}_{i=1}^{n}\cup\{v_{n+j}\}\big{)}\,\Big{|}\,[{\mathcal{D}}_{j}]=[z_{1},\dots,z_{n+j}]\Big{)}\geq 1-\alpha.$ (7)
This leads to (6) by the tower property. Therefore, inverting the criterion on the right-hand side of (6) gives a valid prediction set for $Y_{n+j}$.
The reason why vanilla SCP may fail to achieve selection-conditional coverage like (3) is that, conditional on the selection event, the data points $\{Z_{1},\dots,Z_{n},Z_{n+j}\}$ are no longer exchangeable. In other words, in (7), it is unclear how $V_{n+j}$ is distributed over $v_{1},\dots,v_{n+j}$ if we additionally condition on the selection event. As such, a natural remedy is to find a subset of calibration data that are “exchangeable” with respect to the test point conditioning on the selection event, and leverage their scores to calibrate the prediction of the test unit. We introduce our methods below.
3.2 Conformal inference via a reference set
Fix a test unit $j\in[m]$. Recall that $Z_{n+j}(y)=(X_{n+j},y)$ is the imputed test point with a hypothesized response $y\in{\mathcal{Y}}$. The core of our method is to find calibration points that are exchangeable with respect to the test point conditional on the selection set. At a high level, they are calibration units $i\in[n]$ that are “indistinguishable” with $j\in[m]$, in the sense that treating $Z_{n+j}(y)$ as a calibration point and $Z_{i}$ as a test point results in the same selection event.
We formalize this idea via a “swap” operation. For any calibration unit $i\in[n]$, we define the “swapped” calibration data ${\mathcal{D}}_{{\textnormal{calib}}}^{{\textnormal{swap}(i,j)}}(y)$ and the swapped test data ${\mathcal{D}}_{{\textnormal{test}}}^{{\textnormal{swap}(i,j)}}$ as follows: